OpenStudy (anonymous):

write 2.96 as a mixed number or fraction.

OpenStudy (igreen):

2.96 96 is the decimal. It can be written as 96/100. Your fraction will be 2 96/100 but you need to simplify 96/100.

OpenStudy (igreen):

96/100 can be re-written and simplified as 24/25. So your mixed number will be \(2 \dfrac{24}{25}\).

OpenStudy (igreen):

Or you can write it as an improper fraction by multiplying the denominator by the whole number, and then adding 24. Remember, you keep the denominator. 25 * 2 = 50 50 + 24 = 74 So we have 74/25.